OpenStudy (anonymous):

Samantha and Jake are training for a marathon.In the first week of training Jake ran an average of 7.75 minutes per mile. Later, in week four of training he ran an average of 7.00 minutes per mile. In the first week of training Samantha ran an average of 8.10 minutes per mile. Later, in week six of training she ran an average of 6.60 minutes per mile. In which week will Jake and Samantha have the same average minutes per mile? If Jake and Samantha continue to train until week 33, what will their times be

OpenStudy (anonymous):

in week 8

OpenStudy (anonymous):

how did you figure that out @britbrat4290

OpenStudy (anonymous):

X + 4Y = 32 3X + 10Y = 84 (-16/-2 , -12/-2)........(8 , 6).....My own method I teach my students Mult by 5.......5X + 20Y = 160........20Y = 160 - 5X Mult by 2.......6X + 20Y = 168........20Y = 168 - 6X 160 - 5X = 168 - 6X..........Add 6X 160 + X = 168...................Subtract 160 X = 8 ANSWER in week 8

OpenStudy (anonymous):

what about what there times will be at week 33 is there an equation i can do to figure it out?

OpenStudy (anonymous):

plug it in for x

OpenStudy (anonymous):

plug what in for x?

OpenStudy (anonymous):

the answer u were trying to find in the begggininng 8

OpenStudy (anonymous):

but plug it into what

OpenStudy (anonymous):

X!!!!!!!!!!!!!